Diffusion Morning Report05/18/05

Ted Berg tberg at snf.stanford.edu
Wed May 18 06:31:26 PDT 2005

Good Morning All,
    Looking pretty good this morning. A couple of furnaces have display 
problems-checking it out. Tylan 2 source controller is not cooling.
 Ebeam is in regen. HBr gas cab testing today-gas should be on after 
that. Everything else is OK fore now.

  	Date: 	5/18/2005 	THICKNESS-Change at
Tube Name/Number 	Status 	Comments 	 
Tylan 1 	Up 	  	 
Tylan 2 	yellow 	source controller issues 	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	Up 	  	 
Tylan 8-SiGe 	Up 	  	3.12/30
Tylan 9-Poly 	Up 	  	6.69/30
Tylan 10-Nitride 	yellow 	Observing particles 	11.66/30
Tylan 11-BPSG-LTO 	Up 	  	17.01/30
Tylan 12-FGA 	Up 	  	 
Tystar 	Up 	  	24.78/30
Doped Poly 	GONE 	  	 
TEOS 	yellow 	  	 
Tylan 16 	N/A 	  	 
Thermco 1 	Up 	  	 
Thermco 2 	Up 	  	 
Thermco Poly 	yellow 	testing doped 	 
Thermco Nitride 	Down 	Waiting 	 
Burn Box 	Up 	  	 
ZYGO 	up 	  	.
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
SiH4 for furnaces 	  	B on line 400 PSI 	 changeover working well 

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20050518/7cd1f4c1/attachment.html>

More information about the diffusion mailing list